### v2

# Assessing and Mitigating Medical Knowledge Drift and Conflicts in Large Language Models

Assessing and Mitigating Medical Knowledge Drift and Conflicts in Large Language Models
Weiyi Wu1, Xinwen Xu2, Chongyang Gao3, Xingjian Diao4, Siting Li1, Lucas A. Salas5, Jiang Gui1,4+ {affiliations}
Department of Biomedical Data Science, Dartmouth College, Hanover, NH
Medical Practice Evaluation Center, Massachusetts General Hospital, Boston, MA
Department of Computer Science, Northwestern University, Evanston, IL
Department of Computer Science, Dartmouth College, Hanover, NH
Department of Epidemiology, Dartmouth College, Hanover, NH
$\boldsymbol{+}$ Lead author
1 Abstract
Large Language Models (LLMs) have great potential in the field of health care, yet they face great challenges in adapting to rapidly evolving medical knowledge. This can lead to outdated or contradictory treatment suggestions. This study investigated how LLMs respond to evolving clinical guidelines, focusing on concept drift and internal inconsistencies. We developed the DriftMedQA benchmark to simulate guideline evolution and assessed the temporal reliability of various LLMs. Our evaluation of seven state-of-the-art models across 4,290 scenarios demonstrated difficulties in rejecting outdated recommendations and frequently endorsing conflicting guidance. Additionally, we explored two mitigation strategies: Retrieval-Augmented Generation and preference fine-tuning via Direct Preference Optimization. While each method improved model performance, their combination led to the most consistent and reliable results. These findings underscore the need to improve LLM robustness to temporal shifts to ensure more dependable applications in clinical practice. The dataset is available at https://huggingface.co/datasets/RDBH/DriftMed .
2 Introduction
Healthcare delivery is increasingly strained by the expansion of biomedical knowledge, driven by advancements in research technologies, molecular investigations, drug development, and clinical trials [1, 2, 3]. This proliferation challenges healthcare systems and practitioners, as clinicians face difficulty in staying current, particularly as their formal training may become outdated. Standard practices taught during medical education can soon become obsolete or counterproductive [4, 5], exacerbating the gap between medical education and current best practices. Empirical studies show that clinical guidelines often require reassessment within a few years [6], underscoring the need for innovative methods to support timely and reliable clinical decision-making.
Large language models (LLMs) have emerged as prospective tools for addressing this information proliferation. Recent LLMs, such as GPT‑4o [7] and Llama-3.3 [8], demonstrate physician‑level proficiency in standardized medical exams and can analyze extensive medical literature [9, 10, 11]. Healthcare systems have begun integrating these tools into clinical workflows to facilitate documentation, patient communication, and diagnostic reasoning [12]. However, despite their potential, LLMs present notable limitations. Recent studies on medical LLMs have identified biases in clinical recommendations associated with social determinants of health (SDoH) [13, 14, 15] and cognitive factors [16]. Despite growing attention to SDoH and cognitive factors in LLMs, the temporal factor associated with evolving medical knowledge remains underexplored.
The temporal progression of medical knowledge presents a fundamental challenge for LLMs in healthcare, referred to as concept drift, where established medical guidance changes over time due to the evolution of care standards. Diagnostic criteria, treatment guidelines, and management methods are regularly updated as new evidence emerges. For example, diagnostic criteria for insulin overbasalization (titration of basal insulin beyond an appropriate dose) have shifted from fixed dosage thresholds to contextual markers, including glucose trends and hypoglycemia risk [17]. Similarly, treatment guidelines for HIV/HCV co-infection have evolved from ribavirin-based regimens to newer, safer, and more effective alternatives [18].
Concept drift affects LLMs in two critical dimensions: external concept drift post-deployment and internal knowledge conflict during training.
External Medical Concept Drift: Following deployment, concept drift presents a persistent challenge for LLMs as the gap between static parametric knowledge and the changing medical reality continuously widens. As medical advancements surpass a model’s fixed understanding, concept drift gradually erodes model trustworthiness. Static models developed on the fixed dataset cannot adapt to emerging health threats or incorporate new findings after training, leading to growing misalignment with current clinical practice [19]. This divergence becomes obvious during public health emergencies, such as the COVID-19 pandemic, when the infection control protocols underwent frequent and major revisions, and sticking to outdated recommendations posed increased risks [20, 21].
Internal Medical Knowledge Conflict: During training, LLMs face challenges processing heterogeneous and inconsistently curated data from different periods without the ability to distinguish temporal context [22, 23]. Research demonstrates that LLMs have intrinsic difficulties with training corpora: temporal misalignment, embedded disinformation, and inherent biases [24]. This risk is particularly acute in medicine, where evolving standards are common. LLMs may treat outdated and contemporary recommendations without mechanisms to resolve temporal inconsistencies as equally valid, potentially leading to inappropriate or harmful suggestions. This challenge is further compounded when clinical guidelines undergo reversals over time [25, 26]. For example, the NICE-SUGAR study demonstrated that intensive glucose control in critically ill patients, once thought beneficial based on earlier evidence, actually increased mortality compared to conventional control [25]. Such guideline reversals underscore the urgent need for LLMs not only to incorporate updated evidence but also to unlearn outdated standards that have been shown to be ineffective or dangerous.
Figure 1: Overview of DriftMedQA benchmark construction and prompt example. (Left) Up-to-date clinical guidelines are paired with manually constructed pseudo-outdated counterparts. Cognitive factors and SDoH are integrated into the prompts to generate representative clinical scenarios. (Right) Example advice pairs showing raw guideline content used in the scenario construction. (Bottom) Example of a final model evaluation prompt containing a contextual narrative with embedded self-diagnosis bias.
To systematically assess how internal medical knowledge conflicts and external medical concept drift affect LLM performance and reliability, we developed DriftMedQA, a benchmark designed to simulate the temporal evolution of clinical knowledge. DriftMedQA pairs up-to-date clinical recommendations with manually constructed pseudo-outdated counterparts, enabling direct evaluation of model alignment with evolving standards. To better reflect real-world clinical decision-making complexity, we employed LLMs to generate diverse clinical scenarios based on medical advice, optionally incorporating cognitive factors or SDoH. These enriched contexts simulate realistic patient presentations. A schematic overview of the benchmark construction process and example prompt is provided in Figure 1.
Using this enriched framework, we evaluated seven cutting-edge LLMs across a diverse set of 4,290 annotated clinical scenarios. We provided every LLM with the same prompt to control for variability and ensure consistency for direct comparison, enabling us to quantify how temporal knowledge drift and internal knowledge conflicts affect their reliability. Our baseline findings demonstrate that LLMs consistently struggle to detect when medical knowledge has changed, often failing to embrace up-to-date guidance or to flag flawed, outdated recommendations. Models also exhibited systematic vulnerability to internal conflicts, particularly when encountering subtle but clinically critical changes, where both the old and new recommendations could be misinterpreted as equally current.
To begin addressing these challenges, we explored two complementary mitigation strategies. Retrieval-Augmented Generation (RAG) was employed to incorporate up-to-date external knowledge at inference time, and preference fine-tuning (using Direct Preference Optimization, DPO) was used for parameter-efficient fine-tuning to prioritize correct contemporary advice over outdated standards. While RAG and DPO individually improved model performance, their combined application yielded the greatest benefits, significantly enhancing alignment with current guidelines and reducing internal contradictions. Nevertheless, even with these mitigations, LLMs remained susceptible to reasoning errors in complex clinical environments. Our findings highlight the critical need for continuous updating mechanisms, robust temporal evaluation frameworks, and hybrid strategies that combine external retrieval with internal adaptation to support LLMs’ safe and reliable use in dynamic healthcare settings.
Our work contributes to the growing body of evidence on how knowledge conflicts arise in LLMs and explores preliminary methods for dynamically updating models in response to evolving evidence. Although more robust mechanisms for resolving knowledge drift and reconciling conflicting information are still needed, these findings offer a crucial step toward safer and more effective LLM deployment in healthcare. This work underscores the importance of ongoing efforts to develop dynamic, temporally aware LLM systems by illuminating the pitfalls of static knowledge and knowledge conflicts in an ever-changing field.
3 Results
3.1 Model evaluation
We first assessed the baseline performance of LLMs on the DriftMedQA benchmark. Seven state-of-the-art LLMs were evaluated across 4,290 synthetic clinical scenarios designed to test temporal reliability. To capture diverse patterns of guideline evolution, in the initial evaluation phase, model performance was analyzed across five clinically meaningful update dimensions: Clinical Context, Diagnostic Thresholds, Implementation Approach, Recommendation Intensity, and Treatment Modality. Performance was assessed using three complementary metrics: endorsement of current clinical advice (ECDAadh), rejection of outdated advice (ECDArej), and overall alignment score (ECDAall); see the Methods section for details.
(a) ECDAadh
(b) ECDArej
(c) ECDAall
Figure 2: Evaluation of external medical concept drift. Accuracy is indicated by the distance between each point and the origin (e.g., a radius of 0.9 corresponds to 90% accuracy). Each axis represents a type of modification to clinical guidelines.
All assessed models exhibited varying performance across the five types of clinical recommendation updates (Figure 2). GPT-4o and Qwen2.5-72B demonstrated the highest endorsement scores (ECDAadh), with sample-weighted averages of 0.90 and 0.92, respectively. These scores were significantly higher than the third-best performing model, Qwen2.5-7B (both $p<0.0001$). However, both models exhibited substantial declines when assessed on their ability to reject pseudo-outdated recommendations (ECDArej), with scores of 0.395 for GPT-4o and 0.278 for Qwen2.5-72B. Conversely, Ministral-8B achieved the highest rejection performance (ECDArej $=0.80$), following by gemma-2-27B (ECDArej $=0.68$) and Llama-3-8B (ECDArej $=0.63$), outperforming models with stronger endorsement capabilities (Figure 2(b)).
When considering overall alignment across both current and outdated scenarios, GPT-4o achieved the highest score (ECDAall $=0.65$), as shown in Figure 2(c). This performance was significantly higher than that of the second-best model, Llama-3.3-70B (ECDAall $=0.61$, $p=0.00033$), and the third-best model, Qwen2.5-72B (ECDAall $=0.60$, $p=0.0006$).
Figure 3: Internal medical knowledge conflict across clinical change types. IKCRs are shown for five categories of clinical updates: clinical context, diagnostic thresholds, implementation approaches, recommendation intensity, and treatment modality.
As shown in Figure 3, all evaluated models exhibited measurable internal conflicts, with considerable variability across models and types of guideline updates. Our analysis revealed that more capable or larger-scale models did not consistently exhibit lower internal knowledge conflict ratios (IKCRs; see the Methods section for details). For instance, the 72B parameter version of Qwen2.5 demonstrated higher IKCRs than its 7B counterpart across most evaluated categories. Similarly, Llama-3.3-70B did not show lower conflict ratios compared to Llama-3-8B. Among all models evaluated, the Mistral-8B model achieved the lowest overall IKCR, with a weighted average score of 0.34 across all scenario types, followed by Gemma-2-27B at 0.39.
All evaluated models exhibited knowledge conflicts across all five modification categories. The highest average IKCRs were observed for changes under the groups Implementation Approach (average IKCR $=0.55$) and Treatment Modality (average IKCR $=0.53$). While our baseline evaluation distinguished performance across five guideline change categories, the mitigation analysis focuses on overall alignment and conflict rates to emphasize aggregate improvements.
3.2 Mitigation strategies
We evaluated two complementary strategies to improve temporal reliability and internal consistency in LLM outputs: Retrieval Augmented Generation (RAG) and Preference fine-tuning via Direct Preference Optimization (DPO), as illustrated in Figure4. Each method was tested individually and in combination across different LLMs (see the Methods section for details), with model performance assessed on ECDA and IKCR metrics as summarized in Figure 5. While not exhaustive, these initial strategies helped illuminate where and how to improve temporal alignment.
Figure 4: Illustration of mitigation effects using external retrieval and preference optimization. The left (blue) panel shows model inputs: the top row is the baseline scenario, and the second row adds retrieved external knowledge, representing the RAG-augmented input. The right panels show model outputs from Ministral-8B. The orange panel reflects baseline (top) and RAG-only (bottom) responses; the yellow panel shows DPO-only output (top) and the combined RAG + DPO response (bottom). Only the combined approach yields the correct answer aligned with clinical guidelines.
Application of RAG and DPO independently improved the models’ adherence to current medical guidelines (ECDAadh) relative to their baseline performance (Figure 5(a)). The impact of RAG on the models’ ability to reject outdated advice (ECDArej) was variable across models (Figure. 5(b)). While RAG improved rejection performance for some models, it decreased ECDArej for Ministral-8B and Llama-3-8B compared to their respective baselines.
When considering overall alignment (ECDAall; Figure 5(c)), both RAG and DPO individually improved performance, but the combined application of RAG and DPO consistently yielded the highest ECDAall scores across all models. This improvement from the combined approach was consistently greater than the best-performing single method (RAG or DPO alone) for each model.
Analysis of the IKCR (Figure5(d)) showed that DPO alone reduced IKCR across all evaluated models. RAG alone reduced internal contradictions for most models compared to their baseline. However, for Ministral-8B and Llama-3-8B, applying RAG alone increased IKCR (Ministral-8B: 0.40 vs 0.34 baseline; Llama-3-8B: 0.70 vs 0.45 baseline). Notably, combining RAG with DPO resulted in the lowest IKCR for all models, including Ministral-8B (0.10) and Llama-3-8B (0.26), surpassing the reductions achieved by DPO alone.
(a) ECDAadh
(b) ECDArej
(c) ECDAall
(d) IKCR
Figure 5: Effect of mitigation strategies on model alignment and internal consistency. Each line originates from the baseline performance of a given model and shows changes following the application of RAG (blue), DPO (red), or their combination (yellow). Rightward shifts indicate improvement, while leftward shifts reflect performance degradation. Metrics include endorsement of current advice, rejection of outdated advice, overall alignment, and internal knowledge conflict ratio.
4 Discussion
Our study systematically examined the susceptibility of large language models (LLMs) to external medical concept drift and internal knowledge conflicts when processing clinical recommendations. Using the DriftMedQA benchmark, we found that models often struggle to follow current medical guidelines while failing to reject outdated advice.
A key finding is that LLMs consistently struggled to reject outdated clinical recommendations, even when capable of endorsing current standards. For instance, while models like GPT-4o and Qwen2.5-72B achieve high endorsement rates ($\text{ECDA}_{\text{adh}}$), their rejection rates of outdated advice (ECDArej) were markedly lower, revealing an asymmetric temporal reliability. Furthermore, contrary to expectations, increasing model parameter size did not consistently reduce internal knowledge conflicts. Larger models such as Qwen2.5-72B and Llama-3.3-70B exhibited comparable or even higher Internal Knowledge Conflict Ratios (IKCRs) compared to smaller models, with Ministral-8B achieving the lowest conflict rate.
These vulnerabilities raise important concerns for the safe deployment of LLMs in healthcare. Temporal instability risks clinical decision-making, especially in high-stakes domains such as treatment selection or diagnostic criteria interpretation. Notably, the highest rates of internal conflicts were observed in treatment modality recommendations, where endorsement of conflicting recommendations may lead to inappropriate therapies or dosages. This concern highlights the need for increased attention and a robust evaluation framework in the digital medicine community before widespread clinical application.
To begin addressing these problems, we investigated a set of representative mitigation strategies designed to probe the extent and nature of possible improvements. RAG consistently improved adherence to current guidelines (ECDAadh) across all models, but its effect on rejecting outdated information (ECDArej) was inconsistent. Notably, smaller models like Ministral-8B and Llama-3-8B rely solely on RAG, which paradoxically worsens their ability to identify and reject outdated advice. Since the scenario prompts were identical for both baseline and RAG mitigated evaluations, this performance degradation likely reflects interactions between retrieved documents and the model’s internal parametric knowledge. This suggests that while generally reinforcing correct guidance, retrieval may inadvertently encourage affirmation of flawed or outdated standards if subtle invalidating factors are insufficiently emphasized, leading to incorrect endorsement.
In contrast, DPO fine-tuning directly adapted model parameters to prioritize contemporary clinical guidance. While DPO alone improved overall performance, the combined application of RAG and DPO achieved the most substantial gains across all evaluation metrics ($\text{ECDA}_{\text{all}}$, $\text{ECDA}_{\text{adh}}$, $\text{ECDA}_{\text{rej}}$, and IKCR). This synergy was especially evident in reducing internal knowledge conflicts for smaller models. For instance, in Llama-3-8B, combining RAG and DPO reduced IKCR from 0.45 (baseline) to 0.26, outperforming either method alone (DPO alone: 0.43; RAG alone: 0.70). This suggests DPO not only shapes model preferences but may also enhance the integration or prioritization of external information retrieved via RAG, leading to more coherent and stable reasoning. Our results indicate that hybrid strategies combining external retrieval with internal model refinement offer a promising pathway to strengthen LLMs’ temporal reliability in dynamic clinical settings.
At the advice level, both RAG and DPO can theoretically achieve near-perfect accuracy. RAG can retrieve the recommendation verbatim without context, as each query exactly matches its counterpart in the knowledge base. Also, fine-tuned models can memorize that fixed advice set. However, a more subtle challenge appears when the same models are evaluated on synthetic clinical scenarios, as we did in this analysis. Models performing well in evaluating isolated medical advice perform significantly worse in the face of synthetic clinical scenarios incorporating cognitive factors and SDoH. This discrepancy highlights that the accuracy of knowledge itself cannot ensure reliable performance in complex environments. Real-world medical decision-making rarely occurs in simplified, context-free environments, underscoring the importance of evaluating LLMs on isolated knowledge and their ability to reason within multifaceted, dynamic scenarios.
While we only explored two mitigation strategies that are relatively straightforward to implement, do not require elaborate dataset curation, and have reasonable computational costs, our results demonstrate their potential to improve temporal consistency with current clinical guidelines. Due to a lack of access to proprietary model weights and limited computational resources, we could not apply DPO universally across all assessed models. Additionally, our evaluation was limited to synthetic clinical scenarios that may not fully capture the complexity and diversity of clinical practice. Future work should consider using real-world cases abstracted from healthcare workers with varying levels of complexity, common typographical errors, and incomplete information to better test models’ adaptability and generalization capabilities in realistic medical settings.
Future directions should include developing more well-designed datasets for fine-tuning that explicitly incorporate complex clinical contexts, expanding the range of fine-tuning methods, and advocating for broader access to model parameters. Standardized evaluation protocols that capture real-life clinical complexity could enhance comparative assessments across models. These insights emphasize the need for ongoing efforts to develop dynamically adaptable, temporally reliable, and contextually robust language models to safely support clinical decision-making in rapidly evolving healthcare environments.
5 Methods
5.1 Large Language Models
We evaluated seven state-of-the-art LLMs spanning a range of model sizes and architectures: GPT-4o (OpenAI) [7], Llama-3-8B-Instruct and Llama-3.3-70B-Instruct (Meta) [8], Qwen2.5-7B-Instruct and Qwen2.5-72B-Instruct (Alibaba) [27], Gemma-2-27B-it (Google) [28], and Mixtral-8x7B-Instruct-2410 (Mistral AI) [29].
5.2 Dataset Construction
We developed DriftMedQA, a dataset of 195 clinical recommendation pairs covering chronic (n = 66) and infectious diseases (n = 129). Each pair includes current recommendations (as of February 2025) from the most recent HIV and diabetes clinical guidelines [18, 17], alongside manually created, mutually exclusive, pseudo-outdated versions. We derived these pseudo-outdated recommendations using one of five strategies designed to reflect common patterns of knowledge evolution in clinical guideline updates:
  • Clinical Context (N=22, 11.3%): Revisions to the specific patient populations or clinical circumstances to which a recommendation applies (e.g., narrowing or broadening age ranges, changing applicability based on risk status).
    Examples: revising age applicability from “adults aged <60 years” to “adults aged <70 years”; narrowing recommendation from “all patients” to “only high-risk patients”.
  • Diagnostic & Threshold (N=42, 21.5%): Modifications to specific numerical criteria or classifications used in diagnosis or risk stratification (e.g., changing diagnostic thresholds for blood glucose or HbA1c, altering risk score cutoffs).
    Examples: changing the fasting glucose diagnostic threshold from “100–110 mg/dL” to “110–125 mg/dL”; adjusting HbA1c criteria from $\geq$6.5%” to $\geq$7.0%”.
  • Implementation Approach (N=32, 16.4%): Changes in how care is delivered, organized, or monitored, including methods, processes, systems, duration, or frameworks, even if the core treatment or diagnosis remains similar.
    Examples: shifting from “moderate complexity” to “low complexity” management; transitioning from “lifelong monitoring” to a “short-term surveillance”.
  • Recommendation Intensity (N=53, 27.2%): Changes in the strength or certainty of a recommendation while the core action remains the same (e.g., shifting from permissive to directive language, or vice versa).
    Examples: changing recommendation wording from “may consider” to “should recommend”; from “not recommended” to “recommended” for the same action.
  • Treatment Modality (N=46, 24.6%): Changes in the specific medical interventions recommended (e.g., replacing an older drug class with a newer one, shifting from surgical to non-surgical approaches).
    Examples: replacing “metformin” with “GLP-1 receptor agonists”; transitioning from “surgical intervention” to “physical therapy”.
To further evaluate LLM performance under clinically relevant and cognitively diverse conditions, we transformed each medical recommendation into a richly contextualized, scenario-based question-answer (QA) pair. This design was motivated by prior work highlighting the impact of cognitive biases and SDoH on LLM clinical reasoning [16, 13, 15, 14]. Each scenario was conditioned on one of ten cognitive or social factors commonly encountered in medical decision-making, with an additional neutral “No Factor” setting in which no cognitive factor or SDoH was introduced. The selected factors — self‑diagnosis, recency, confirmation, frequency, status quo, cultural, socioeconomic, racial or ethnic, geographic, and false consensus — capture realistic variation in reasoning without introducing factual distortion or adversarial intent. We used Qwen2.5-72B [27] to generate these scenarios by systematically combining each medical recommendation with its corresponding factor. This pipeline produced a total of $4{,}290$ scenario‑based QA pairs $(11\leavevmode\nobreak\ \text{factors}\times 195\leavevmode\nobreak\ \text{recommendation}\times 2\leavevmode\nobreak\ )$, evenly split between current and pseudo‑outdated recommendations.
5.3 Evaluation Metrics
We evaluated LLMs’ temporal reliability through two complementary metrics: one quantifying alignment with external evolving medical guidelines and the other detecting internal knowledge inconsistencies.
[…]
$\text{ECDA}_{\text{adh}}$ $=\frac{1}{|\mathcal{D}_{U}|}\sum_{s_{i,c,u}\in\mathcal{D}_{U}}\mathbf{1}(\hat{y}_{i,c,u}=1)$ (1)
$\text{ECDA}_{\text{rej}}$ $=\frac{1}{|\mathcal{D}_{O}|}\sum_{s_{i,c,o}\in\mathcal{D}_{O}}\mathbf{1}(\hat{y}_{i,c,o}=0)$ (2)
$\text{ECDA}_{\text{all}}$ $=\frac{\text{ECDA}_{\text{adh}}+\text{ECDA}_{\text{rej}}}{2}$ (3)
$\text{ECDA}_{\text{adh}}$ (Eq. 1) measures the model’s ability to correctly endorse current medical guidelines ($y_{i,c,u}=1$), while $\text{ECDA}_{\text{rej}}$ (Eq. 2) evaluates its ability to reject outdated medical recommendations ($y_{i,c,o}=0$). Their average $\text{ECDA}_{\text{all}}$ (Eq. 3) provides a balanced assessment of temporal guideline discrimination.
Internal Knowledge Conflict Ratio (IKCR). To assess internal consistency, we evaluated whether models simultaneously endorsed conflicting recommendations. Our evaluation scenarios present paired current ($s_{i,c,u}$) and outdated ($s_{i,c,o}$) versions for each core clinical concept $i$ and change $c$. Let $\hat{y}_{i,c,u}$ and $\hat{y}_{i,c,o}$ be the model’s binary predictions ($1=\text{endorse}$). We define the set of active pairs, $\mathcal{A}$, as those where the model endorses at least one version ($\mathcal{A}=\{(i,c)\mid\hat{y}_{i,c,u}=1\lor\hat{y}_{i,c,o}=1\}$). An internal contradiction, or knowledge conflict, occurs for an active pair $(i,c)\in\mathcal{A}$ when the model simultaneously endorses both mutually exclusive recommendations ($\hat{y}_{i,c,u}=1\land\hat{y}_{i,c,o}=1$). The IKCR quantifies the frequency of such contradictions:
$\text{IKCR}=\frac{\sum_{(i,c)\in\mathcal{A}}\mathbf{1}(\hat{y}_{i,c,u}=1\land\hat{y}_{i,c,o}=1)}{|\mathcal{A}|}$ (4)
A higher IKCR indicates a greater frequency of internal logical contradictions, which could undermine clinical reliability.
5.4 Mitigating Strategies for Evolving Medical Knowledge
We explored two complementary and orthogonal strategies to address this challenge: non-parametric knowledge augmentation and parametric knowledge adaptation. Non-parametric augmentation was applied to all evaluated LLMs except Qwen2.5-72B. Due to limited training resources and lack of access to proprietary model weights, parametric adaptation and its combination with retrieval was evaluated only on Qwen2.5-7B, Mistral-8B, and Llama-3-8B.
Non-Parametric Knowledge Augmentation: This strategy supplements the model with external information during inference without modifying its internal parameters. Specifically, we employed Retrieval-Augmented Generation (RAG), using a knowledge base of 195 up-to-date clinical advice.
For each clinical query scenario $s$, we encode the query using Sentence-BERT encoders [30, 31] and retrieve the top-$k$ most relevant guideline snippets ($d_{i}$) from our knowledge base ($\mathcal{KB}$) based on cosine similarity, then augment the input prompt with these documents before generating the response:
$D_{k}$ $=\underset{d_{i}\in\mathcal{KB}}{\operatorname{TopK}}\left(\cos\left(E_{q}(\text{query}(s)),E_{d}(d_{i})\right),k\right)$ (5)
$\hat{y}_{s}$ $=\text{LLM}\!\bigl{(}s\oplus D_{k};\theta_{\text{base}}\bigr{)}.$ (6)
where $E_{q}$ and $E_{d}$ are query and document encoders, $\cos$ denotes cosine similarity, $k=2$ in our experiments, $\oplus$ represents concatenation, and $\theta_{\text{base}}$ denotes the unchanged base model parameters. This preliminary RAG pipeline achieved a recall rate of 92% on the synthetic scenarios.
This non‑parametric strategy delivers clear clinical benefits: it decouples the model from its knowledge source, allows guideline updates to be incorporated instantly without retraining, and retains explicit citations to authoritative documents. Those advantages, however, come with costs. The knowledge base demands continual curation and governance; each inference step triggers a retrieval call, adding latency and operational complexity; system performance depends on the coverage and freshness of external sources; and retrieval errors can introduce hallucinations or amplify existing biases.
Parametric Knowledge Adaptation: While non-parametric methods update knowledge outside the model, parametric approaches directly modify the model’s weights. These approaches include supervised fine-tuning (SFT) methods [32] and preference-based approaches leveraging reinforcement learning (RL) [33, 34]. For our investigation, we explored Direct Preference Optimization (DPO) [35], a preference fine-tuning method that avoids the need for an explicit reward model by refining the model through direct comparisons between candidate outputs.
This approach was particularly well-suited for medical knowledge updates because it naturally accommodates our core challenge: simultaneously encouraging adherence to current guidelines while discouraging outdated advice. Unlike SFT, which requires extensively and carefully curated question-answer datasets, DPO operates directly on preference triplets $(x,y_{w},y_{l})$. Here, for a given clinical advice input $x$ (derived from our dataset), $y_{w}$ represents a response indicating endorsement of the correct guideline version (chosen), and $y_{l}$ represents endorsement of the incorrect version (rejected).
The DPO framework effectively implements a form of targeted unlearning by penalizing incorrect clinical advice ($y_{l}$) while reinforcing correct standards ($y_{w}$) through its objective function. The DPO objective and our parameter-efficient implementation fine-tuning approach are defined as:
$\mathcal{L}_{\text{DPO}}(\theta_{\text{new}})$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{\text{pref}}}\left[\log\sigma\left(\beta\log\frac{p_{\theta_{\text{new}}}(y_{w}|x)}{p_{\text{ref}}(y_{w}|x)}-\beta\log\frac{p_{\theta_{\text{new}}}(y_{l}|x)}{p_{\text{ref}}(y_{l}|x)}\right)\right]$ (7)
$\theta_{\text{new}}$ $=\theta_{\text{base}}+\Delta\theta_{\text{LoRA}}$ (8)
where $\mathcal{D}_{\text{pref}}$ is the dataset of preference triplets; $p_{\theta_{\text{new}}}$ is the fine-tuned policy model; $p_{\text{ref}}$ is the reference model (the base LLM before DPO fine-tuning); $\sigma$ is the logistic function; and $\beta$ is a scaling hyperparameter. In Eq. 8 only a small subset of Low-Rank Adaptation (LoRA) [36] parameters $\Delta\theta_{\text{LoRA}}$ are updated while the base parameters $\theta_{\text{base}}$ remain frozen. We configured LoRA with rank $r=8$ and scaling factor $\alpha=16$. DPO fine-tuning was conducted using two NVIDIA A100 80GB GPUs. The process continued until the model achieved 100% accuracy on the pseudo-outdated versus up-to-date advice pairs, thereby ensuring complete memorization of the clinical recommendations. The model was then evaluated on independent synthetic scenarios to assess its ability to generalize this memorized knowledge to unseen clinical contexts.
6 Code & Data availability
The dataset used in this study is publicly available at https://huggingface.co/datasets/RDBH/DriftMed. The code supporting this study will be released upon publication.
7 Acknowledgment
This project is supported by the Department of Defense’s Congressionally Directed Medical Research Programs (DOD CDMRP) Award HT9425-23-1-0267.